# A four-operator splitting algorithm for nonconvex and nonsmooth optimization

### v2

@@ Abstract
In this work, we address a class of nonconvex nonsmooth optimization problems where the objective function is the sum of two smooth functions (one of which is proximable) and two nonsmooth functions (one weakly convex and proximable and the other continuous and weakly concave). We introduce a new splitting algorithm that extends the Davis-Yin splitting (DYS) algorithm to handle such four-term nonconvex nonsmooth problems. We prove that with appropriately chosen step sizes, our algorithm exhibits global subsequential convergence to stationary points with a stationarity measure converging at a rate of $1/k$. When specialized to the setting of the DYS algorithm, our results allow for larger stepsizes compared to existing bounds in the literature. Experimental results demonstrate the practical applicability and effectiveness of our proposed algorithm.
@@ 1 Introduction
The functions, $f$, $g$, $h$ and $p$ satisfy the following:
@@
    $g:{\rm I}\!{\rm R}^{n}\to(-\infty,\infty]$ is a proper, closed, and $\rho_{{}_{\rm{g}}}$-weakly convex function; that is, $g+\frac{\rho_{{}_{\rm{g}}}}{2}\|\cdot\|^{2}$ is convex where $\rho_{{}_{\rm{g}}}\geq 0$;
@@
    $p:{\rm I}\!{\rm R}^{n}\to{\rm I}\!{\rm R}$ is continuous (possibly nonsmooth) on an open set containing the domain of $g$ such that $-p$ is $L_{{}_{\rm{p}}}$-weakly convex with $L_{{}_{\rm{p}}}\geq 0$; that is, $\frac{L_{{}_{\rm{p}}}}{2}\|\cdot\|^{2}-p$ is convex.
A notable special case of (1.1) is the three-term optimization problem, where $p\equiv 0$. The Davis-Yin splitting (DYS) algorithm, introduced in (Davis and Yin, 2017), tackles this problem for convex functions $f$, $g$, and $h$, assuming $h$ additionally meets Assumption 1.1(c). To extend the algorithm’s applicability to nonconvex problems, Bian and Zhang (2021) leveraged techniques from (Li and Pong, 2016) to show that with a sufficiently small stepsize, the DYS algorithm achieves global subsequential convergence. This requires $f$ and $h$ to satisfy conditions (a) and (c) of Assumption 1.1, respectively, and $g$ to be any proper closed function. However, despite not requiring the functions to be convex, this extension is still limited as it can only admit at most one nonsmooth function.
@@
DC optimization.
The DYS algorithm is not applicable to difference-of-convex (DC) regularized optimization problems, which take the form
$\min_{x\in{\rm I}\!{\rm R}^{n}}~{}f(x)+g(x)-\hat{g}(x),$ (1.2)
where $g$ and $\hat{g}$ are convex. The limitation arises from the potential nonsmoothness of the concave component $-\hat{g}$. Meanwhile, by setting $h\equiv 0$ and $p\coloneqq-\hat{g}$, we satisfy Assumption 1.1 (c) and (d) with $L_{{}_{\rm{h}}}=L_{{}_{\rm{p}}}=0$.
@@
By noting the convexity of the function on the right-hand side of (1.4) (being the supremum of affine (convex) functions), we may then fit (1.3) in the form (1.1) by setting $h\equiv 0$ and $p\coloneqq\sum_{i=1}^{r}M_{\lambda_{i}P_{i}}\circ A_{i}$ so that Assumption 1.1(d) holds with $L_{{}_{\rm{p}}}=\sum_{i=1}^{r}\lambda_{i}^{-1}$.
@@
We can also draw examples from feasibility problems reformulated as optimization problems: Let $A,B,C$ and $D_{i}$ for $i\in\{1,2,\dots,r\}$ be closed nonempty sets where $A$, $B$, $C$ are convex, and the $D_{i}$’s are not necessarily convex. The feasibility problem is given by:
@@
$\min_{x\in{\rm I}\!{\rm R}^{n}}~{}\underbrace{\frac{1}{2}d^{2}_{A}(x)}_{f(x)}+\underbrace{\frac{1}{2}d^{2}_{B}(x)}_{g(x)}+\underbrace{\frac{1}{2}d^{2}_{C}(x)}_{h(x)}+\underbrace{\sum_{i=1}^{r}\frac{1}{2}d^{2}_{D_{i}}(x)}_{p(x)}.$ (1.5)
@@
The original setting of the DYS algorithm in (Davis and Yin, 2017) can handle this problem when $r=0$. On the other hand, the extended setting considered in (Bian and Zhang, 2021) can handle the case $B={\rm I}\!{\rm R}^{n}$ and $r=1$ (i.e., the feasibility problem with two convex sets and one nonconvex set), but it requires a very small stepsize. Meanwhile, by the convexity of $A$, $B$ and $C$, we see that Assumption 1.1(a), (b) and (c) hold for Eq. 1.5. The function $p$, on the other hand, is a nonsmooth function due to the nonconvexity of each $D_{i}$. Nevertheless, noting that the half-squared distance function is the Moreau envelope of the indicator function, we have from (1.4) that $p$ satisfies Assumption 1.1(d) with $L_{{}_{\rm{p}}}=r$.
In this paper, we propose a splitting algorithm that generalizes the DYS algorithm to the four-term optimization problem (1.1), which encompasses a broader range of nonconvex and nonsmooth problem classes, including those described above. We prove that our algorithm is globally subsequentially convergent to stationary points for appropriately chosen stepsizes, with a first-order optimality measure converging at a rate of $1/k$. Another major contribution of this work is the derivation of upper bounds for stepsizes for such convergence results. When specialized to $p\equiv 0$, our results yield stepsize estimates for the DYS algorithm that are significantly larger than the existing bounds in (Bian and Zhang, 2021). Up to our knowledge, for this special case of $p\equiv 0$, our iteration complexity result is also new for the DYS algorithm. During the final stages of preparing this manuscript, we came across a recent preprint by Dao et al. (2024) that examines a specific case of our problem in which $L_{{}_{\rm{p}}}=0$ in Assumption 1.1(d). Notably, our analysis yields upper bounds for stepsizes that are considerably larger than those reported in their work.
@@ 3.1 Proposed algorithm
When $L_{{}_{\rm{p}}}=0$, we in particular set $\beta=\infty$, so that $\gamma=\alpha$ by Step 0 of Algorithm 1. On the other hand, when $L_{{}_{\rm{f}}}=L_{{}_{\rm{h}}}=0$, we set $\alpha=\infty$ and therefore $\gamma=\beta$.
@@
    $y^{k+1}={\rm prox}_{\gamma g}(y^{k}-\gamma\partial p(y^{k}),$
@@
    (Proximal DC algorithm). When $f\equiv 0$, the algorithm simplifies to
@@
    Thus, when $\tau=1$ and $L_{{}_{\rm{p}}}=0$ so that $\beta=\infty$ and $\gamma=\alpha$, we further obtain
    $y^{k+1}\in{\rm prox}_{\gamma g}(y^{k}-\gamma\nabla h(y^{k})-\gamma\partial\phi(y^{k})),$ (3.6)
    which is the proximal DC algorithm (Wen et al., 2018) for solving (1.2).
@@ 3.2 Stepsize bounds
We now derive appropriate stepsizes for Algorithm 1 that will guarantee sufficient descent of some merit function. First, we introduce some notations. Given $\xi\in\partial p(y)$, we define $V_{\Lambda,\xi}:{\rm I}\!{\rm R}^{n}\times{\rm I}\!{\rm R}^{n}\times{\rm I}\!{\rm R}^{n}\to{\rm I}\!{\rm R}$ as
@@
Suppose that $f$ and $h$ are continuously differentiable functions, $g$ satisfies Assumption 1.1 (b), and $\gamma>0$ satisfies $\frac{1}{\gamma}=\frac{1}{\alpha}+\frac{1}{\beta}$ and $\gamma\leq\frac{1}{\rho_{{}_{\rm{g}}}}$, with $\alpha,\beta>0$. Then $P_{\Lambda}(y,z)\neq\emptyset$ and
@@
Given $y^{\scriptscriptstyle+}\in\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\Phi_{\Lambda,\xi}(w;y,z,x)$ for some $x\in{\rm prox}_{\alpha f}(z)$ and $\xi\in\partial p(y)$, we have from the optimality condition of (3.7) that
$0\in\nabla f(x)+\nabla h(x)+\frac{1}{\alpha}(y^{\scriptscriptstyle+}-x)+\left(\xi+\frac{1}{\beta}(y^{\scriptscriptstyle+}-y)\right)+\partial g(y^{\scriptscriptstyle+}).$ (3.10)
Rearranging the terms, we see that
00 0 $\in\left(\frac{1}{\alpha}+\frac{1}{\beta}\right)y^{\scriptscriptstyle+}-\frac{1}{\alpha}\left(x-\alpha\nabla f(x)-\alpha\nabla h(x)\right)-\frac{1}{\beta}\left(y-\beta\xi\right)+\partial g(y^{\scriptscriptstyle+})$ (3.11)
Multiplying the above by $\gamma$ and noting that $\frac{1}{\gamma}=\frac{1}{\alpha}+\frac{1}{\beta}$, we obtain
$0\in y^{\scriptscriptstyle+}-\frac{\gamma}{\alpha}T_{\alpha(f+h)}(x)-\frac{\gamma}{\beta}T_{\beta p}(y)+\gamma\partial g(y^{\scriptscriptstyle+}).$ (3.12)
Meanwhile, we note that
[…]
Since $g$ is $\rho_{{}_{\rm{g}}}$-weakly convex, $\gamma\leq\frac{1}{\rho_{{}_{\rm{g}}}}$, and (3.12) holds, we conclude that $y^{\scriptscriptstyle+}$ is an element of the set on the left-hand side of (3.13) by looking at its optimality condition. Consequently, $y^{\scriptscriptstyle+}\in P_{\Lambda}(y,z)$ by (3.13) and therefore $\supseteq$” in (3.9) holds. The other inclusion can be proved by reversing the above arguments and noting that our setting for $\gamma,\alpha,\beta$ ensures that $\Phi_{\Lambda,\xi}(w;y,z,x)$ is convex with respect to $w$, so Eq. 3.10 is also a sufficient condition for optimality. ∎
@@
$V_{\Lambda,\xi}(y,z,x)\leq Q_{\alpha(f+h)}(y;x)+p(y)-\left(\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}\right){\left\|{y-y^{\scriptscriptstyle+}}\right\|}^{2}+g(y)$
@@
Given $\xi\in\partial p(y)$, note that
$Q_{\beta p}(y^{\scriptscriptstyle+};y,\xi)=p(y)+\left\langle\xi,y^{\scriptscriptstyle+}-y\right\rangle+\frac{1}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}.$ (3.14)
Meanwhile,
$Q_{\alpha f}(y^{\scriptscriptstyle+};x)$ $=f(x)+\left\langle\nabla f(x),y^{\scriptscriptstyle+}-x\right\rangle+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-x}\right\|}^{2}$
$=\left(f(x)+\left\langle\nabla f(x),y-x\right\rangle+\frac{1}{2\alpha}{\left\|{y-x}\right\|}^{2}\right)+\left\langle\nabla f(x),y^{\scriptscriptstyle+}-y\right\rangle$
$\qquad+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}+\frac{1}{\alpha}\left\langle y-x,y^{\scriptscriptstyle+}-y\right\rangle$
$=Q_{\alpha f}(y;x)+\frac{1}{\alpha}\left\langle y-T_{\alpha f}(x),{y^{\scriptscriptstyle+}-y}\right\rangle+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}.$
It follows that
$Q_{\alpha(f+h)}(y^{\scriptscriptstyle+};x)=Q_{\alpha(f+h)}(y;x)+\frac{1}{\alpha}\left\langle y-T_{\alpha(f+h)}(x),{y^{\scriptscriptstyle+}-y}\right\rangle+\frac{1}{2\alpha}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}.$ (3.15)
Using (3.14) and (3.15) and noting (3.8), we have
$\begin{split}\Phi_{\Lambda,\xi}(y^{\scriptscriptstyle+};y,z,x)&=Q_{\alpha(f+h)}(y;x)+p(y)+\frac{1}{2\gamma}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}\\ &\quad+\left\langle\frac{1}{\alpha}\left(y-T_{\alpha(f+h)}(x)\right)+\xi,y^{\scriptscriptstyle+}-y\right\rangle+g(y^{\scriptscriptstyle+})\end{split}$ (3.16)
On the other hand, since $y^{\scriptscriptstyle+}\in\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}\Phi_{\Lambda,\xi}(w;y,z,x)$, we have from (3.11) and the $\rho_{{}_{\rm{g}}}$-weak convexity of $g$ that
$g(y)$ $\geq g(y^{\scriptscriptstyle+})+\left\langle\frac{1}{\alpha}T_{\alpha(f+h)}(x)+\frac{1}{\beta}(y-\beta\xi)-\frac{1}{\gamma}y^{\scriptscriptstyle+},y-y^{\scriptscriptstyle+}\right\rangle-\frac{\rho_{{}_{\rm{g}}}}{2}{\left\|{y-y^{\scriptscriptstyle+}}\right\|}^{2}$
$=g(y^{\scriptscriptstyle+})+\left\langle\frac{1}{\alpha}T_{\alpha(f+h)}(x)+\frac{1}{\beta}(y-\beta\xi)-\frac{1}{\gamma}y,y-y^{\scriptscriptstyle+}\right\rangle-\left(\frac{\rho_{{}_{\rm{g}}}}{2}-\frac{1}{\gamma}\right){\left\|{y-y^{\scriptscriptstyle+}}\right\|}^{2}.$ (3.17)
Combining (3.16) and (3.17), we obtain
$\Phi_{\Lambda,\xi}(y^{\scriptscriptstyle+};y,z,x)$ $\leq Q_{\alpha(f+h)}(y;x)+p(y)+g(y)+\left(\frac{1}{2\gamma}-\frac{1}{\gamma}+\frac{\rho_{{}_{\rm{g}}}}{2}\right){\left\|{y-y^{\scriptscriptstyle+}}\right\|}^{2}.$
Simplifying the last expression and using Eq. 3.7, we obtain the claim of the lemma. ∎
@@
Suppose that Assumption 1.1 holds. Let $\rho_{{}_{\rm{f}}}\geq 0$ be such that $f$ is $\rho$-weakly convex and $L\geq L_{{}_{\rm{f}}}$ be such that $L-\rho_{{}_{\rm{f}}}>0$. Then
$\begin{split}&V_{\Lambda,\xi}(y,z,x)\\ \geq&Q_{\alpha(f+h)}(y^{\scriptscriptstyle+};\hat{x})+p(y^{\scriptscriptstyle+})+g(y^{\scriptscriptstyle+})+\left\langle\nabla f(\hat{x})-\nabla f(x)-\frac{1}{\alpha}(\hat{x}-x),x-y^{\scriptscriptstyle+}\right\rangle\\ &\quad+\left\langle\nabla h(x)-\nabla h(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}\\ &\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{\hat{x}-x}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}\end{split}$ (3.18)
@@
$\partial\left(\frac{L_{{}_{\rm{p}}}}{2}\|\cdot\|^{2}-p\right)(y)=L_{{}_{\rm{p}}}y-\partial p(y),$ (3.21)
@@
$\frac{L_{{}_{\rm{p}}}}{2}{\left\|{y^{\scriptscriptstyle+}}\right\|}^{2}-p(y^{\scriptscriptstyle+})\geq\frac{L_{{}_{\rm{p}}}}{2}{\left\|{y}\right\|}^{2}-p(y)+\left\langle L_{{}_{\rm{p}}}y-\xi,y^{\scriptscriptstyle+}-y\right\rangle,$
@@
$p(y)\geq p(y^{\scriptscriptstyle+})-\left\langle\xi,y^{\scriptscriptstyle+}-y\right\rangle-\frac{L_{{}_{\rm{p}}}}{2}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}$ (3.22)
@@
$Q_{\beta p}(y^{\scriptscriptstyle+};y,\xi)\geq\left(p(y^{\scriptscriptstyle+})+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}\right).$ (3.23)
@@
Regarding the requirement of $\rho_{{}_{\rm{f}}}$ in Lemma 3.4, by Assumption 1.1(a) and Remark 2.2, we see that there indeed exists $\rho_{{}_{\rm{f}}}\geq 0$ such that $f+\frac{\rho_{{}_{\rm{f}}}}{2}{\left\|{\cdot}\right\|}^{2}$ is convex.
@@
$V_{k-1}-V_{k}$ $\geq-\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta L_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}+\frac{1-\gamma\rho_{{}_{\rm{g}}}}{2\gamma}{\left\|{y^{k}-y^{k+1}}\right\|}^{2}$ (3.24)

### v4

@@ Abstract
In this work, we address a class of nonconvex nonsmooth optimization problems where the objective function is the sum of two smooth functions (one of which is proximable) and two nonsmooth functions (one proper, closed and proximable, and the other continuous and weakly concave). We introduce a new splitting algorithm that extends the Davis-Yin splitting (DYS) algorithm to handle such four-term nonconvex nonsmooth problems. We prove that with appropriately chosen stepsizes, our algorithm exhibits global subsequential convergence to stationary points with a stationarity measure converging at a global rate of $1/T$, where $T$ is the number of iterations. When specialized to the setting of the DYS algorithm, our results allow for larger stepsizes compared to existing bounds in the literature. Experimental results demonstrate the practical applicability and effectiveness of our proposed algorithm.
@@ 1 Introduction
The functions, $f$, $g$, $h$ and $p$ satisfy the following: [^1] See Section 2 for the definitions of $L$-smooth, $\rho_{g}$-weakly convex, and proximal mapping.
@@
    $g:{\rm I}\!{\rm R}^{n}\to(-\infty,\infty]$ is a proper, closed, and proximable function;
@@
    $p:{\rm I}\!{\rm R}^{n}\to{\rm I}\!{\rm R}$ is continuous (possibly nonsmooth) on an open set containing the domain of $g$ such that $-p$ is $\rho_{{}_{\rm{p}}}$-weakly convex with $\rho_{{}_{\rm{p}}}\geq 0$.
A notable special case of (1.1) is the three-term optimization problem, where $p\equiv 0$. The Davis-Yin splitting (DYS) algorithm, introduced in (Davis and Yin, 2017), tackles this problem for convex functions $f$, $g$, and $h$, assuming $h$ additionally meets Assumption 1.1(c). To extend the algorithm’s applicability to nonconvex problems, Bian and Zhang (2021) leveraged techniques from (Li and Pong, 2016) to show that with a sufficiently small stepsize, the DYS algorithm achieves global subsequential convergence. This requires $f$, $g$ and $h$ to satisfy conditions (a), (b) and (c) of Assumption 1.1, respectively. However, despite not requiring the functions to be convex, this extension of the DYS algorithm is still limited as it can only admit at most one nonsmooth function.
@@
Nonsmooth nonconvex optimization with concave component.
The DYS algorithm is not applicable to problems that take the form
$\min_{x\in{\rm I}\!{\rm R}^{n}}~{}F(x)+G(x)-H(x),$ (1.2)
where $F$ is $L$-smooth, $G$ is proper and closed, and $H$ is convex. The limitation of the DYS algorithm arises from the potential nonsmoothness of the concave component $-H$. Meanwhile, by setting $f=F$, $g=G$, $h\equiv 0$ and $p\coloneqq-H$ in (1.1), we satisfy Assumption 1.1 with $L_{{}_{\rm{h}}}=\rho_{{}_{\rm{p}}}=0$.
@@
By noting the convexity of the function on the right-hand side of (1.4) (being the supremum of affine (convex) functions), we may then fit (1.3) in the form (1.1) by setting $f\equiv 0$ and $p\coloneqq\sum_{i=1}^{r}M_{\lambda_{i}P_{i}}\circ A_{i}$ so that Assumption 1.1(d) holds with $\rho_{{}_{\rm{p}}}=\sum_{i=1}^{r}\lambda_{i}^{-1}$.
@@
We can also draw examples from feasibility problems reformulated as optimization problems: Let $A,B,C$ and $D_{i}$ for $i\in\{1,2,\dots,r\}$ be closed nonempty sets where $A$ and $C$ are convex, and the sets $B$ and $D_{i}$ $(i\in\{1,\dots,r\})$ are not necessarily convex. The feasibility problem is given by:
@@
$\min_{x\in{\rm I}\!{\rm R}^{n}}~{}\underbrace{\frac{1}{2}d^{2}_{A}(x)}_{f(x)}+\underbrace{\delta_{B}(x)}_{g(x)}+\underbrace{\frac{1}{2}d^{2}_{C}(x)}_{h(x)}+\underbrace{\sum_{i=1}^{r}\frac{1}{2}d^{2}_{D_{i}}(x)}_{p(x)}.$ (1.5)
@@
The original setting of the DYS algorithm in (Davis and Yin, 2017) can handle this problem when $r=0$. On the other hand, the extended setting considered in (Bian and Zhang, 2021) can handle the case $B={\rm I}\!{\rm R}^{n}$ and $r=1$ (i.e., the feasibility problem with two convex sets and one nonconvex set), but it requires a very small stepsize. Meanwhile, by the convexity of $A$, $B$ and $C$, we see that Assumption 1.1(a), (b) and (c) hold for Eq. 1.5. The function $p$, on the other hand, is a nonsmooth function due to the nonconvexity of each $D_{i}$. Nevertheless, noting that the half-squared distance function is the Moreau envelope of the indicator function, we have from (1.4) that $p$ satisfies Assumption 1.1(d) with $\rho_{{}_{\rm{p}}}=r$.
We note that the optimization problem (1.1) under Assumption 1.1 can be reformulated into the form (1.2) by setting $H\coloneqq p-\frac{\rho_{{}_{\rm{p}}}}{2}{\left\|{\cdot}\right\|}^{2}$ and choosing either (i) $F\coloneqq f+h+\frac{\rho_{{}_{\rm{p}}}}{2}{\left\|{\cdot}\right\|}^{2}$, $G\coloneqq g$, or (ii) $F\coloneqq f+h$, $G\coloneqq g+\frac{\rho_{{}_{\rm{p}}}}{2}{\left\|{\cdot}\right\|}^{2}$. Under such reformulations, one may apply existing algorithms such as the generalized proximal point (GPP) method An and Nam (2016), which coincides with the proximal difference-of-convex (PDC) algorithm Wen et al. (2018) when $F$ and $G$ are convex.
However, one of the advantages of the framework we propose is that it enables a direct treatment of the problem (1.1) without requiring such reformulations, and it takes advantage of the decomposition of the smooth term of the objective into $f+h$, where $f$ and $h$ possess the structures specified in Assumption 1.1(a) and (c), respectively. [^2] In practice, when both terms satisfy Assumption 1.1(a), it is often more practical to assign more structure to $f$, as the proximal step tends to exploit more information than the gradient step.
1.1 Contributions
In this paper, we propose an extension of the Davis–Yin splitting (DYS) algorithm to handle optimization problems involving two nonsmooth terms, thereby expanding its application to a broader class of nonconvex and nonsmooth problems, including those described earlier. Specifically, we develop a four-operator splitting algorithm for the general problem (1.1) under the setting described in Assumption 1.1. Our main contributions are summarized below:
  1. (I)
    We prove that our algorithm, which includes the DYS splitting as a special case, is globally subsequentially convergent to stationary points under appropriately chosen stepsizes. Furthermore, a first-order optimality measure is shown to converge at a global rate of $1/T$, where $T$ is the number of iterations.
  2. (II)
    A key theoretical contribution of this work is the derivation of upper bounds on the stepsizes that guarantee convergence. When specialized to the case of $p\equiv 0$, our results yield stepsize estimates for the DYS algorithm that are significantly larger than existing bounds in (Bian and Zhang, 2021). As a practical consequence, our experiments show that these larger stepsizes lead to faster convergence of the algorithm.
  3. (III)
    Our experiments demonstrate that leveraging the structures of $f$ and $h$ separately leads to faster convergence compared to GPP/PDC, which treats the smooth term as a single component. Hence, while reformulation into (1.2) is possible, our direct approach offers practical advantages due to utilization of individual structures.
During the final stage of preparing the first draft of this paper, we came across the preprint Dao et al. (2024), which introduces the “doubly relaxed forward-Douglas–Rachford (DRFDR)” splitting algorithm. This method applies to a special case of our problem setting where $\rho_{{}_{\rm{p}}}=0$ in Assumption 1.1(d). When $\rho_{{}_{\rm{p}}}>0$, we note that DRFDR can also be applied to (1.1) by subtracting $\frac{\rho_{{}_{\rm{p}}}}{2}{\left\|{\cdot}\right\|}^{2}$ from $p$, and then adding $\frac{\rho_{{}_{\rm{p}}}}{2}{\left\|{\cdot}\right\|}^{2}$ to either $f$, $g$ or $h$. However, this transformation introduces an additional step, and determining where it is most practical to add this quadratic term is not straightforward. In contrast, our proposed algorithm accommodates the case $\rho_{{}_{\rm{p}}}>0$ without requiring such transformations. More importantly, our convergence analysis relies on entirely different proof techniques, enabling us to derive larger stepsize bounds compared to those in Dao et al. (2024). Our numerical experiments demonstrate that these larger stepsizes lead to improved practical performance, highlighting another key distinction between our work and Dao et al. (2024).
1.2 Organization
@@ 3.1 Proposed algorithm
When $\rho_{{}_{\rm{p}}}=0$, we in particular set $\beta=\infty$, so that $\gamma=\alpha$ by Step 0 of Algorithm 1. On the other hand, when $L_{{}_{\rm{f}}}=L_{{}_{\rm{h}}}=0$, we set $\alpha=\infty$ and therefore $\gamma=\beta$.
@@
    $y^{k+1}={\rm prox}_{\gamma g}(y^{k}-\gamma\partial p(y^{k})),$
@@
    (Generalized proximal point algorithm/Proximal DC algorithm). When $f\equiv 0$, the algorithm simplifies to
@@
    Thus, when $\tau=1$ and $\rho_{{}_{\rm{p}}}=0$ so that $\beta=\infty$ and $\gamma=\alpha$, we further obtain
    $y^{k+1}\in{\rm prox}_{\gamma g}(y^{k}-\gamma\nabla h(y^{k})-\gamma\partial p(y^{k})).$ (3.6)
    When $g$ is convex, this method is known as the proximal DC (PDC) algorithm (Wen et al., 2018), while for a general (not necessarily convex) $g$, it is referred to as the generalized proximal point (GPP) algorithm in An and Nam (2016).
@@ 3.2 Stepsize bounds
We now derive appropriate stepsizes for Algorithm 1 that will guarantee sufficient descent of a certain merit function. First, we introduce some notations. Given $\xi\in\partial p(y)$, we define $V_{\Lambda,\xi}:{\rm I}\!{\rm R}^{n}\times{\rm I}\!{\rm R}^{n}\times{\rm I}\!{\rm R}^{n}\to{\rm I}\!{\rm R}$ as
@@
Suppose that $f$ and $h$ are continuously differentiable functions, $g$ is proper and closed, and $\gamma>0$ satisfies $\frac{1}{\gamma}=\frac{1}{\alpha}+\frac{1}{\beta}$ with $\alpha,\beta>0$. Then $P_{\Lambda}(y,z)\neq\emptyset$ and
@@
Given $x\in{\rm prox}_{\alpha f}(z)$ and $\xi\in\partial p(y)$, we have that
$~{}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}~{}\Phi_{\Lambda,\xi}(w;y,z,x)$
$=$ $~{}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}~{}f(x)+h(x)+\left\langle\nabla f(x)+\nabla h(x),w-x\right\rangle+\frac{1}{2\alpha}{\left\|{w-x}\right\|}^{2}$
$\qquad\qquad\quad+p(y)+\left\langle\xi,w-y\right\rangle+\frac{1}{2\beta}{\left\|{w-y}\right\|}^{2}+g(w)$
$=$ $~{}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}~{}\left\langle\nabla f(x)+\nabla h(x),w\right\rangle+\frac{1}{2\alpha}{\left\|{w}\right\|}^{2}-\frac{1}{\alpha}\left\langle w,x\right\rangle$
$\qquad\qquad\quad+\left\langle\xi,w\right\rangle+\frac{1}{2\beta}{\left\|{w}\right\|}^{2}-\frac{1}{\beta}\left\langle w,y\right\rangle+g(w)$
$=$ $~{}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}~{}\frac{1}{2\gamma}{\left\|{w}\right\|}^{2}-\frac{1}{\alpha}\left\langle x-\alpha(\nabla f(x)+\nabla h(x)),w\right\rangle-\frac{1}{\beta}\left\langle y-\beta\xi,w\right\rangle+g(w)$
$=$ $~{}\operatorname*{arg\,min}_{w\in{\rm I}\!{\rm R}^{n}}~{}\frac{1}{2\gamma}{\left\|{w-\left(\frac{\gamma}{\alpha}(x-\alpha\nabla f(x)-\alpha\nabla h(x))+\frac{\gamma}{\beta}(y-\beta\xi)\right)}\right\|}^{2}+g(w)$
$=$ $~{}{\rm prox}_{\gamma g}\left(\frac{\gamma}{\alpha}(x-\alpha\nabla f(x)-\alpha\nabla h(x))+\frac{\gamma}{\beta}(y-\beta\xi)\right)$
where the third equality follows from the definition of $\gamma$. The claim now follows from (3.5). ∎
@@
$V_{\Lambda,\xi}(y,z,x)\leq Q_{\alpha(f+h)}(y;x)+p(y)+g(y)$
@@
We immediately get the result by noting from (3.7) that $V_{\Lambda,\xi}(y,z,x)\leq\Phi_{\Lambda,\xi}(y;y,z,x)$. ∎
@@
[…]
$\begin{split}&V_{\Lambda,\xi}(y,z,x)\\ \geq&Q_{\alpha(f+h)}(y^{\scriptscriptstyle+};\hat{x})+p(y^{\scriptscriptstyle+})+g(y^{\scriptscriptstyle+})+\left\langle\nabla f(\hat{x})-\nabla f(x)-\frac{1}{\alpha}(\hat{x}-x),x-y^{\scriptscriptstyle+}\right\rangle\\ &\quad+\left\langle\nabla h(x)-\nabla h(\hat{x}),y^{\scriptscriptstyle+}-\hat{x}\right\rangle+\frac{1}{2(L-\rho_{{}_{\rm{f}}})}{\left\|{\nabla f(x)-\nabla f(\hat{x})}\right\|}^{2}\\ &\quad-\left(\frac{L_{{}_{\rm{h}}}}{2}+\frac{1}{2\alpha}+\frac{\rho_{{}_{\rm{f}}}L}{2(L-\rho_{{}_{\rm{f}}})}\right){\left\|{\hat{x}-x}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}\end{split}$ (3.10)
@@
$\partial\left(\frac{\rho_{{}_{\rm{p}}}}{2}\|\cdot\|^{2}-p\right)(y)=\rho_{{}_{\rm{p}}}y-\partial p(y),$ (3.13)
@@
$\frac{\rho_{{}_{\rm{p}}}}{2}{\left\|{y^{\scriptscriptstyle+}}\right\|}^{2}-p(y^{\scriptscriptstyle+})\geq\frac{\rho_{{}_{\rm{p}}}}{2}{\left\|{y}\right\|}^{2}-p(y)+\left\langle\rho_{{}_{\rm{p}}}y-\xi,y^{\scriptscriptstyle+}-y\right\rangle,$
@@
$p(y)\geq p(y^{\scriptscriptstyle+})-\left\langle\xi,y^{\scriptscriptstyle+}-y\right\rangle-\frac{\rho_{{}_{\rm{p}}}}{2}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}$ (3.14)
@@
$Q_{\beta p}(y^{\scriptscriptstyle+};y,\xi)\geq\left(p(y^{\scriptscriptstyle+})+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{\scriptscriptstyle+}-y}\right\|}^{2}\right).$ (3.15)
@@
$V_{k-1}-V_{k}$ $\geq-\frac{c(\alpha)}{2\tau\alpha}{\left\|{x^{k}-x^{k-1}}\right\|}^{2}+\frac{1-\beta\rho_{{}_{\rm{p}}}}{2\beta}{\left\|{y^{k}-y^{k-1}}\right\|}^{2}$ (3.16)